OpenStudy (anonymous):

In a direct variation, y is 4 when x is 12. What is y when x = 6?

OpenStudy (anonymous):

I think the answer is 2 but I'm not so sure so Im asking for a check up

OpenStudy (anonymous):

yes the answer is two

OpenStudy (fifciol):

y=kx so 4=12k so k must be 1/3 so y=1/3*6=2

OpenStudy (anonymous):

because to get x from 12 to 6 you divide by two so you also have to divide the y (4) by two so your answer is then two

OpenStudy (fifciol):

your answer is correct

OpenStudy (anonymous):

@Fifciol @ses11 thank you both